# Embedding in Recommender Systems: A Survey

### v2

@@ 5.4. Survey and Tool
Tools: EasyRec  [^13] https://github.com/alibaba/EasyRec provides a convenient tool for users to develop and customize the recommendation model. The novel part of EasyRec is that it integrates the autoML API from Alibaba Cloud, which supplies lots of autoML-related services including auto feature selection, and auto feature interaction.
5.5. Future Direction
Progressive NAS in EDS: Progressive NAS (PNAS) is a novel neural architecture search (NAS) scheme that leverages the shortest path algorithm to discover the most efficient path between a start and target node within the search space. Considering the extensive search space involved in embedding size, PNAS holds significant promise for its application in the search for optimal embedding sizes in the field of EDS.
@@ 5.5. Future Direction
Unsupervised NAS in embedding: Conventionally, the best model architecture is determined based on human-labeled data. However, (Liu et al., 2020a) proposed a new unsupervised scheme of NAS called UnNAS in visual tasks. The performance of the sampled architecture needs to be evaluated in an unsupervised manner. In addition, the model selected by UnNAS has comparable performance with that chosen by the supervised NAS method. Thus, it is a promising pattern in NAS where we could explore an efficient method of EDS.
@@ 6. Quantization
Quantization, a fundamental technique in information theory, emerges as a promising avenue to bolster the scalability of deep recommender systems. It aims to compress continuous high-dimensional vectors into low-dimensional discrete codes. In deep recommender systems, there can be tens of thousands of categorical features that are encoded into dense embeddings. The huge memory cost caused by the embeddings poses a challenge to the deployment of deep recommender systems. In this context, quantization offers a strategic pathway to alleviate the memory burden associated with embeddings, thereby contributing to the enhanced scalability and operational efficiency of recommendation systems.
Quantization compresses the original embedding into a set of codes. The length of this set of codes will be much smaller than the original embedding. One can either (1) use such discrete codes as new embeddings or (2) reconstruct the original embedding with a small distortion. Therefore, quantization allows us to discard the original embedding, leading to pronounced reductions in memory consumption and offering an effective way to address the scalability challenges in recommender systems.
In this section, we classify quantization methods into: (a) binary quantization and (b) codebook quantization based on the form of quantization embedding. In particular, we additionally discuss the application of quantization embedding on online recommender systems.
Figure 6. Comparison of PQ-based and AQ-based methods
@@ 6.1. Binary Quantization
In particular, to obtain more compact and high-quality codes, DPR adds the balance constraint and the irrelevant constraint. The balance constraint encourages maximizing the entropy of the code, and the irrelevant constraint encourages the bits in the code to be as independent as possible. The objective function of DPR can be formulated as the following:
(23) $\min$ $\sum_{(u,i,j)\in D_{S}}\frac{1}{|U|\left|I_{u}^{+}\right|\left|I_{u}^{-}\right|}\left(2r-\mathbf{b}_{u}^{T}\left(\mathbf{d}_{i}-\mathbf{d}_{j}\right)\right)^{2}$
s.t. $\mathbf{B}\in\{\pm 1\}^{r\times n},\mathbf{D}\in\{\pm 1\}^{r\times m}$
$\mathbf{B}\mathbf{1}_{n}$ $=0,\mathbf{D}\mathbf{1}_{m}=0,\mathbf{BB}^{T}=n\mathbf{I}_{r},\mathbf{D}\mathbf{D}^{T}=m\mathbf{I}_{r}$
Where $\mathbf{B}$ and $\mathbf{D}$ denote the matrices of $b$ and $d$ stacked by columns. For this discrete optimization problem, DPR solves it with softening constraints and alternating optimization. However, DPR does not impose a constraint on the gap between the binary code and the initial embedding, which is improved by Discrete Deep Learning (DDL) (Zhang et al., 2018). DDL uses a bag-of-words model to learn the embedding of the text of an item and minimizes the gap between it and the corresponding binary code.
@@ 6.2. Codebook Quantization
Codebook quantization is a family of methods that compress the original embedding matrix into a set of codebooks and indexes so that we can reconstruct the embedding that will eventually be used for downstream recommendation tasks with codebooks and indexes that only require a small memory footprint. Here, we present three types of codebook quantization methods: unsupervised and supervised codebook quantization.
@@
[…]
@@
Vanilla PQ does not optimize the objective directly, it applies K-means clustering to each of the $M$ subspaces and compresses each subspace into $K$ codewords, so that each feature vector is compressed into quantization embedding $i(\mathbf{x})=\{1,...,K\}^{M}$. By computing the Euclidean distance, Vanilla PQ can be applied to the nearest neighbor search. However, Vanilla PQ has a serious problem in that it directly divides the features into different subspaces with equal spacing in an ordered manner and may ignore the possibility of a high correlation between subspaces, which can substantially degrade the performance of quantization. Optimized Product Quantization (OPQ) (Ge et al., 2013) uses the rotation matrix $R\in\mathbb{R}^{M\times M}$ to optimize the decomposition of subspaces to reduce the correlation between subspaces, so the objective function is modified as:
(25) $\min_{R,\mathcal{C}^{1},\ldots,\mathcal{C}^{M}}\sum_{\mathbf{x}}\|\mathbf{x}-\mathbf{c}(i(\mathbf{x}))\|^{2}$
[…]
Since the above objective function is difficult to optimize, OPQ adopts the strategy of fixing $R$ and codebook $C$ one at a time to optimize them alternatively.
@@
Both PQ and AQ-based quantization methods work in an unsupervised manner, employing unsupervised k-means clustering or heuristic beam search algorithms to optimize their objective functions. However, these methods fail to address the distortion of embeddings caused by quantization, leading to a partial loss of original embedding information. Consequently, obtaining high-quality embedding becomes challenging. Differentiable Product Quantization (DPQ)  (Chen et al., 2020b) proposes softmax-based and centroid-based methods to minimize reconstruction loss approximately between raw embeddings and quantization embeddings. This approach leads to higher-quality quantization embeddings compared to unsupervised PQ algorithms.
Supervised methods (Chen et al., 2020b; Yue et al., 2016) enhance quantization via reconstruction loss, but recent research (Zhan et al., 2021; Xiao et al., 2021a; Lian et al., 2020a) shows it’s insufficient. Tailoring objectives for specific tasks in recommender systems proves more effective. Product Quantized Collaborative Filtering (PQCF) (Lian et al., 2020b) challenges separate user and item quantization using PQ. Misaligned coordinates make this suboptimal. PQCF minimizes rating prediction loss, moving from Euclidean to the inner product space. It rotates ${\mathbf{C}_{U}}$ and ${\mathbf{C}_{V}}$ with orthogonal matrix $\mathbf{H}$, aligning spaces and addressing PQ-based methods issues. The rating estimated by PQCF can be written as:
(27) $\hat{r}_{ij}=\left(\mathbf{H}^{T}{\mathbf{C}_{U}}\mathbf{b}_{i}\right)^{T}\left(\mathbf{H}^{T}{\mathbf{C}_{V}}\mathbf{d}_{j}\right)$
where $\mathbf{b}_{u}$, $\mathbf{d}_{i}$ denote the quantization embedding of user $i$ and item $j$, respectively. PQCF adds regularization to the rating prediction loss to form the objective function:
(28) $\min_{\mathbf{C}_{U},\mathbf{C}_{V},B,D}\sum_{i,j}w_{ij}\Big{(}r_{ij}$ $-\hat{r}_{ij}\Big{)}^{2}+\lambda\Big{(}\left\|{\mathbf{C}_{U}}\right\|_{F}^{2}+\left\|{\mathbf{C}_{V}}\right\|_{F}^{2}\Big{)}$
Alongside optimizing downstream recommendations, new optimization goals have emerged. Distill-VQ (Xiao et al., 2022), inspired by knowledge distillation, sets its objective as a similarity function measuring differences in relevance score distributions between teacher and student models (e.g., KL divergence). Another innovation, LightRec (Lian et al., 2020a), enhances reconstruction loss with two functions, minimizing differences in user-item ratings pre- and post-quantization, as well as alterations in recommendation ranking. Matching-oriented Product Quantization (MoPQ) (Xiao et al., 2021a) shows that better quantization reconstruction doesn’t always mean better downstream performance. MoPQ improves accuracy through contrastive learning, modeling query-quantization matching via multinoulli process.
@@ 6.3. Online Quantization
Modern recommender systems face constant influxes of new users, and traditional quantization methods (PQ, AQ) lack the capacity to manage streaming data, thus presenting challenges in real-world applications. Similarly, prior online hash algorithms require recalculating all user embeddings upon new user arrivals, which is often impractical in recommendation systems. To tackle this, Online PQ (Xu et al., 2018) posits that the impact of new data on the codebook is minimal due to the smaller size of incremental data. It updates only codewords for new data points without altering quantization embedding for original data. Online PQ maintains a sliding window for data processing, continuously applying K-means clustering for new codewords. However, being unsupervised, Online PQ doesn’t optimize quantization error, risking information loss and low-quality embedding. To address this, Online OPQ (Liu et al., 2020b) extends to streaming data by solving the orthogonal procrustes problem to ensure subspace orthogonality during codebook updates. Online AQ (Liu et al., 2021e) extends AQ, maintaining consistent objective functions. For streaming data adaptability, Online AQ derives codebook update strategies and related regret bounds via linear regression closed solutions and matrix inversion lemma (Tylavsky and Sohie, 1986). Unlike AQ’s beam search for codeword selection, Online AQ introduces an efficient block beam search, a trade-off between hill climbing and beam search.
Finally, we summarize all the quantization algorithms mentioned in this section in Table 5, where you can find the objective function, training method, and whether the algorithm is feasible for online quantization, etc.
Table 5. Summary of Quantization Embedding Methods. QD=Quantization Distortion; RPL=Rating Prediction Loss; BPR=Bayesian Personalized Ranking Loss; BCE=Binary CrossEntropy; MCL=Multinoulli Contrastive Loss; KDL=Knowledge Distillation Loss.
Type Method Codebook size Supervised/Unsupervised Objective function
Binary Quantization DPR (Zhang et al., 2017) No codebook Supervised AUC averaged
DDL (Zhang et al., 2018) No codebook Supervised QD+RPL
Codebook Qantization PQ (Jegou et al., 2010), OPQ (Ge et al., 2013) D/M Unsupervised QD
AQ (Babenko and Lempitsky, 2014) D Unsupervised QD
DPQ (Chen et al., 2020b) D/M Supervised QD
PQCF (Lian et al., 2020b) D/M Supervised QD+RPL
LightRec (Lian et al., 2020a) D/M Supervised QD+BPR+KDL
xLightFM (Jiang et al., 2021) Learnable Supervised QD+BCE
MOPQ (Xiao et al., 2021a) D/M Unsupervised QD+MCL
Distill-VQ (Xiao et al., 2022) D/M Unsupervised QD+KDL
Online Qantization Online PQ (Xu et al., 2018) D/M Unsupervised QD
Online OPQ (Liu et al., 2020b) D/M Unsupervised QD
Online AQ (Liu et al., 2021e) D Unsupervised QD
6.4. Survey and Tool
@@
Survey To the best of our knowledge, no survey on quantization embedding for recommender systems has been conducted prior to this. A concurrent study (Li et al., 2023a) during the same period discussed vector quantization in the context of compressing embeddings in recommender systems. However, their primary focus was on the precision of embedding compression, whereas our emphasis lies in discussing the applications of quantized embeddings across various scenarios within recommender systems, including quantization embedding learning in online recommendation contexts, along with specific quantization techniques. Moreover, existing surveys on quantization either concentrate on image coding (Nasrabadi and King, 1988) or are restricted to product quantization (Ali and Engliš, 2005). Some existing reviews on codebooks also belong to the scope of quantization. For example, (Lu and Chang, 2010) discusses a variety of VQ Codebook Generation algorithms, and (Ramanan and Niranjan, 2012) discusses codebook design in the field of object recognition.
Tool NEQ (Dai et al., 2019) provides various implementations of quantization algorithms (e.g. PQ, AQ, and OPQ). We also refer the reader to a vector quantization repository [^14] https://github.com/lucidrains/vector-quantize-pytorch that implements some novel quantization algorithms(e.g. Residual VQ (Lee et al., 2022)).
6.5. Future Direction
@@
Multi-task Quantization Previous studies (Zhan et al., 2021; Xiao et al., 2021a; Lian et al., 2020a) have highlighted that optimizing solely for reconstruction loss is inadequate for producing high-quality quantization embeddings. For better quantization, tailoring learning tasks to specific downstream recommendation contexts can enhance pattern capture. Introducing multi-task learning objectives alongside quantization distortion minimization as a new objective empowers models to grasp more valuable patterns, potentially leading to improved generalization.

### v3

@@ 6.4. Surveys and Future Directions
Survey. In terms of Deep Recommender Systems (DRS), Zheng et al. (Zheng et al., 2022) and Chen et al. (Chen et al., 2022b) both demonstrate how to perform the autoML technique on each component of DRS, including feature selection, feature embedding, and feature interaction with a different taxonomy of methods. Tools. EasyRec [^5] https://github.com/alibaba/EasyRec provides a convenient tool for users to develop and customize the recommendation model. The novel part of EasyRec is that it integrates the autoML API from Alibaba Cloud, which supplies lots of autoML-related services, including auto feature selection and auto feature interaction.
@@
Future Direction: Unsupervised NAS in embedding. Conventionally, the best model architecture is determined based on human-labeled data. However, (Liu et al., 2020a) proposed a new unsupervised scheme of NAS called UnNAS in visual tasks. The performance of the sampled architecture needs to be evaluated in an unsupervised manner. In addition, the model selected by UnNAS has comparable performance with that chosen by the supervised NAS method. Thus, it is a promising pattern in NAS where we could explore an efficient method.
@@ 7. Quantization
Quantization, a fundamental technique in information theory, emerges as a promising avenue to bolster the scalability of deep recommender systems. It aims to compress continuous high-dimensional vectors into low-dimensional discrete codes. In deep recommender systems, there can be tens of thousands of categorical features that are encoded into dense vectors, also known as embeddings. The high memory cost associated with embeddings poses a challenge to the deployment of deep recommender systems. In this context, quantization provides a strategic pathway to alleviate the memory burden associated with embeddings, thereby enhancing the scalability and operational efficiency of recommender systems.
Quantization compresses the original embedding into a set of codes. The memory cost of this set of codes is much smaller than the original embedding. One can either (1) use such discrete codes as new embeddings or (2) reconstruct the original embedding with a small distortion. Therefore, quantization allows us to discard the original embedding, leading to pronounced reductions in memory consumption and offering an effective way to address the scalability challenges in recommender systems. In this section, we classify quantization methods into: (a) binary quantization and (b) codebook quantization based on the form of quantization embedding. In particular, we additionally discuss the application of quantization embedding on online recommender systems. These quantization approaches are summarized in Table 7. Fig. 7 provides a comprehensive overview of these quantization approaches.
Table 7. Embedding Quantization
Quantization Binary Quantization (§7.1) Key Models: Discrete Personalized Ranking (DPR) (Zhang et al., 2017), Discrete Deep Learning (DDL) (Zhang et al., 2018)
Codebook Quantization (§7.2) Unsupervised Methods: Product Quantization (PQ) (Jegou et al., 2010), Optimized PQ (OPQ) (Ge et al., 2013), Additive Quantization (AQ) (Babenko and Lempitsky, 2014)
Supervised Methods: Differentiable PQ (DPQ) (Chen et al., 2020b), Product Quantized CF (PQCF) (Lian et al., 2020b), LightRec (Lian et al., 2020a), Distill-VQ (Xiao et al., 2022), MoPQ (Xiao et al., 2021a), xLightFM (Jiang et al., 2021)
Online Quantization (§7.3) Key Models: Online PQ (Xu et al., 2018b), Online OPQ (Liu et al., 2020b), Online AQ (Liu et al., 2021f)
Figure 7. A taxonomy of quantization methods discussed in this survey. (Left) Binary Quantization converts embeddings into binary codes. (Center) Codebook Quantization, such as Product Quantization, compresses vectors into a set of discrete indices from codebooks. (Right) Online Quantization adapts codebooks to streaming data.
@@ 7.1. Binary Quantization
In particular, to obtain more compact and high-quality codes, DPR adds the balance constraint and the irrelevant constraint. The balance constraint encourages maximizing the entropy of the code, and the irrelevant constraint encourages the bits in the code to be as independent as possible. For this discrete optimization problem, DPR solves it with softening constraints and alternating optimization. However, DPR does not impose a constraint on the gap between the binary code and the initial embedding, which is improved by Discrete Deep Learning (DDL) (Zhang et al., 2018). DDL uses a bag-of-words model to learn the embedding of the text of an item and minimizes the gap between it and the corresponding binary code.
@@ 7.2. Codebook Quantization
Codebook quantization is a family of methods that compress the original embedding matrix into a set of codebooks and indices, allowing for the reconstruction of the embedding used for downstream recommendation tasks. These codebooks and indexes require a small memory footprint, making them suitable for use in various applications. Here, we present three types of codebook quantization methods: unsupervised and supervised codebook quantization.
@@
$\text{ s.t. }\quad\mathbf{c}\in\mathcal{C}=\mathcal{C}^{1}\times\cdots\times\mathcal{C}^{M}$
@@
Vanilla PQ does not optimize the objective directly; it applies K-means clustering to each of the $M$ subspaces and compresses each subspace into $K$ codewords, so that each feature vector is compressed into a quantization embedding $i(\mathbf{x})=\{1,...,K\}^{M}$. By computing the Euclidean distance, Vanilla PQ can be applied to the nearest neighbor search. However, Vanilla PQ has a serious problem in that it directly divides the features into different subspaces with equal spacing in an ordered manner and may ignore the possibility of a high correlation between subspaces, which can substantially degrade the performance of quantization. Optimized Product Quantization (OPQ) (Ge et al., 2013) uses the rotation matrix $R\in\mathbb{R}^{M\times M}$ to optimize the decomposition of subspaces to reduce the correlation between subspaces.
@@
Both PQ and AQ-based quantization methods work in an unsupervised manner, employing unsupervised k-means clustering or heuristic beam search algorithms to optimize their objective functions. However, these methods fail to address the distortion of embeddings caused by quantization, leading to a partial loss of original embedding information. Consequently, obtaining high-quality embedding becomes challenging. Differentiable Product Quantization (DPQ) (Chen et al., 2020b) proposes softmax-based and centroid-based methods to minimize reconstruction loss approximately between raw embeddings and quantization embeddings. This approach consistently leads to notably higher-quality quantization embeddings compared to traditional unsupervised PQ algorithms.
Supervised methods (Chen et al., 2020b; Yue et al., 2016) enhance quantization via reconstruction loss, but recent research (Zhan et al., 2021; Xiao et al., 2021a; Lian et al., 2020a) shows it’s insufficient. Tailoring objectives for specific tasks in recommender systems proves more effective. Product Quantized Collaborative Filtering (PQCF) (Lian et al., 2020b) challenges separate user and item quantization using PQ. Misaligned coordinates make this suboptimal. PQCF minimizes rating prediction loss, moving from Euclidean to the inner product space. It rotates $\mathbf{C}_{U}$} and $\mathbf{C}_{V}$} with orthogonal matrix $\mathbf{H}$, aligning spaces and addressing PQ-based methods issues. Alongside optimizing downstream recommendations, new optimization goals have emerged. Distill-VQ (Xiao et al., 2022), inspired by knowledge distillation, sets its objective as a similarity function measuring differences in relevance score distributions between teacher and student models (e.g., KL divergence). Another innovation, LightRec (Lian et al., 2020a), enhances reconstruction loss with two functions, minimizing differences in user-item ratings pre- and post-quantization, as well as alterations in recommendation ranking. Matching-oriented Product Quantization (MoPQ) (Xiao et al., 2021a) shows that better quantization reconstruction doesn’t always mean better downstream performance. MoPQ improves accuracy through contrastive learning, modeling query-quantization matching via multinoulli process.
@@ 7.3. Online Quantization
Modern recommender systems face constant influxes of new users, and traditional quantization methods (PQ, AQ) lack the capacity to manage streaming data, thus presenting challenges in real-world applications. Similarly, prior online hash algorithms require recalculating all user embeddings upon new user arrivals, which is often impractical in recommendation systems. To tackle this, Online PQ (Xu et al., 2018b) posits that the impact of new data on the codebook is minimal due to the smaller size of incremental data. It updates only codewords for new data points without altering quantization embedding for original data. Online PQ maintains a sliding window for data processing, continuously applying K-means clustering for new codewords. However, being unsupervised, Online PQ doesn’t optimize quantization error, risking information loss and low-quality embedding. To address this, Online OPQ (Liu et al., 2020b) extends to streaming data by solving the orthogonal procrustes problem to ensure subspace orthogonality during codebook updates. Online AQ (Liu et al., 2021f) extends AQ, maintaining consistent objective functions. For streaming data adaptability, Online AQ derives codebook update strategies and related regret bounds via linear regression closed solutions and matrix inversion lemma (Tylavsky and Sohie, 1986). Unlike AQ’s beam search for codeword selection, Online AQ introduces an efficient block beam search, a trade-off between hill climbing and beam search.
7.4. Surveys and Future Directions
@@
Survey. To the best of our knowledge, there does not exist any survey on quantization embedding for recommender systems. In addition, existing surveys on quantization either focus on image coding (Nasrabadi and King, 1988) or are limited to product quantization (Ali and Engliš, 2005). Some existing reviews on codebooks also belong to the scope of quantization. For example, (Lu and Chang, 2010) discusses a variety of VQ Codebook Generation algorithms, and (Ramanan and Niranjan, 2012) discusses codebook design in the field of object recognition.
@@
Future Direction 1: Multi-task Quantization. Previous studies (Zhan et al., 2021; Xiao et al., 2021a; Lian et al., 2020a) have highlighted that optimizing solely for reconstruction loss is inadequate for producing high-quality quantization embeddings. For better quantization, tailoring learning tasks to specific downstream recommendation contexts can enhance pattern capture. Introducing multi-task learning objectives alongside quantization distortion minimization as a new objective empowers models to grasp more valuable patterns, potentially leading to improved generalization.